01 March 2013

A fantastic day of measurements

Today was another fantastic day for solar atmospheric observations. The sky was clear, and all our instruments were up and running. The temperature at the Eureka Weather Station remains below -50°C. Yesterday's official minimum temperature was -51.4°C, which is impressively cold! The temperature outside the PEARL Ridge Lab was again warmer than at Eureka, at -38°C mid-day. There was minimal wind.

It was another highly productive day in the PEARL Ridge Lab's IR lab. Dan and Joseph took 71 MIR measurements with the Bruker, while Debora took 44 measurements with PARIS. The suntracker worked smoothly throughout the day.

In the UV-Vis Lab, the PEARL-GBS, SAOZ, and UT-GBS all took measurements. Xiaoyi finished stray light and polarization data analysis for the UT-GBS, which showed that the new fibre bundle solved the polarization issue. Resolution tests showed the new optical alignment he performed yesterday improved the UT-GBS' spectral resolution. Xiaoyi continues to work on limiting the stray light reaching the UT-GBS' detector, and reducing the signal level.

While Xiaoyi has succeeded in reducing the signal level (which had been saturating the detector) by a factor of 20, the UT-GBS' CCD continues to have a relatively high signal for winter conditions. He has reduced the CCD exposure time to 30 milliseconds. This is close to the minimum exposure time of 13 milliseconds, and does not allow sufficient room for adjustment as the Sun becomes more intense over the coming months. While he could limit the CCD area used during summer, it is ideal to keep this consistent year-round. Xiaoyi tested a variety of neutral density filters and a diffuser to decrease signal levels, and is in the process of analyzing the results. He will also test modifications to the telescope optics tomorrow, which may provide another avenue of controlling the signal strength.

Volodya started his day by fixing Brewer #192 at the Eureka Weather Station, which had stopped working. After bringing it inside, he found that one of the micrometers had got dirty from grease residue and became very difficult to rotate. He cleaned it and it works well now. Having the Brewer inside also allowed Volodya to do a UV radiation calibration using 50 W reference lamps. The rest of the day, he worked on making the MAESTRO/SPS scripts more robust, aiming to avoid the need for operators to restart the software often. Both MAESTRO and SPS worked well all day.

This morning at 0PAL, Chris and Emily ran a white light calibration for the polarization system. Emily then analyzed the calibration data and will start depolarization analysis tomorrow. Chris made measurements all day, except during a one hour interruption when the PEARL internet connection went down.

In other good news, Zen fixed the E-AERI hatch. His extensive tests of the analogue rain sensor showed it is responsive and accurate. With the hatch problem solved, the E-AERI is now operating nominally, and has been taking data continuously since Wednesday afternoon. Zen has requested information about replacing the digital rain sensor from ABB, and awaits their reply. He also performed standard maintenance checks, and is planning to install a remote power bar to enable the ability to remotely restart the E-AERI computer.

Today's ACE ozonesonde flight occurred at 18:32 (local time) and reached an altitude of 25,980 m. The Raven balloon failed to launch properly due to a tear in the balloon fabric. Fortunately, the Eureka Weather Station SAO (James Allen) quickly stepped on the cord between the sonde and the balloon, cut the cord, and re-launched the sonde on a regular balloon.

Let's all hope the perfect weather continues throughout the campaign!

Dan

Back to Daily Reports
Back to Home